# A Wireless Foundation Model for Multi-Task Prediction

### v1

@@ III-B Model Architecture
Casual Transformer Backbone. We use a casual Transformer to extract features from the input embeddings, consisting of a stack of Transformer layers. Each Transformer layer contains a multi-head causal self-attention mechanism followed by a feed-forward network. For each head $h=\{1,\cdots,H\}$ in the multi-head attention mechanism, we transform $\mathbf{E}^{(m)}$ into a query matrix $\mathbf{Q}^{(m)}_{h}=\mathbf{E}^{(m)}\mathbf{W}^{Q}_{h}$, a key matrix $\mathbf{K}^{(m)}_{h}=\mathbf{E}^{(m)}\mathbf{W}^{K}_{h}$, and a value matrix $\mathbf{V}^{(m)}_{h}=\mathbf{E}^{(m)}\mathbf{W}^{V}_{h}$, where $\mathbf{W}^{Q}_{h},\mathbf{W}^{K}_{h}\in\mathbb{R}^{d_{m}\times d_{k}}$ and $\mathbf{W}^{V}_{h}\in\mathbb{R}^{d_{m}\times d_{v}}$. After that, the $h$-th head after casual attention is given by
@@
where $\mathbf{O}^{(m)}\in\mathbb{R}^{N_{p}\times d_{m}}$ represents the output of multi-head casual attention, and $\mathbf{W}^{O}\in\mathbb{R}^{Hd_{v}\times d_{m}}$ linearly projects these heads to align with the latent dimension of the Transformer. The attention output $\mathbf{O}^{(m)}$ passes through the subsequent feed-forward network and other stacked Transformer layers in the backbone, and is finally converted into a feature matrix $\mathbf{Z}^{(m)}\in\mathbb{R}^{N_{p}\times d_{m}}$.
@@ IV-C1 Baselines and Performance Metric
  • CRLNet: A convolutional long short-term memory (CLSTM) network specifically designed for ISAC scenarios [3]. It combines CNNs for spatial feature extraction with LSTM units to capture temporal dynamics, and is tailored for predictive beamforming via AoA forecasting.
@@ IV-C2 Performance Analysis
We compare the NMSE performance of the foundation models with various baseline schemes across different scenarios, as shown in Fig. 5. As the speeds of the four distinct motion scenarios increase from left to right, a consistent deterioration in performance is observed across all schemes. CRLNet, a model specifically designed for AoA prediction in ISAC, outperforms the naive RNN and Informer networks, both of which were not optimized for ISAC-related tasks. Upon comparing the performance of different schemes, it becomes evident that our proposed foundation models, whether in a single-task or multi-task configuration, significantly outperform the other baseline approaches. Additionally, the single-task model, trained on a specific task, demonstrates a remarkable performance improvement due to its tailored optimization. A more detailed examination of Fig. 5(a) and Fig. 5(b) shows that, in low-speed scenarios (such as Pedestrian and Bicycle), the effect of the prediction interval $\Delta t$ on the performance of all schemes is minimal. In contrast, in high-speed scenarios (such as Vehicle and Rural Vehicle), performance declines as $\Delta t$ increases. This degradation is attributed to the fact that, in high-speed environments, longer prediction intervals result in larger object displacements, which makes it more difficult to predict the AoA accurately. As a result, larger values of $\Delta t$ lead to a more significant reduction in performance.
@@ IV-C3 Downstream Task
Metric RNN CRLNet Informer Foundation Model (Multi-task) Foundation Model (Single-task)
@@ IV-E Zero-Shot and Scaling Analysis
In the previous subsections, we evaluated the zero-shot generalization ability of our foundation model across different configurations within the same task. In this subsection, we take a further step and investigate the model’s zero-shot performance on an unseen task, thereby assessing its capacity for out-of-task generalization. Specifically, we consider a new task: predicting the time-delay in an ISAC system, which corresponds to estimating the physical distance between a user and the BS, i.e. $\nu_{k,t}$ in (8). Structurally, this task resembles previous angle prediction tasks, such as the one described in (II-B). In the following results, we compare the NMSE across time steps for various models to evaluate their performance on the unseen delay prediction task. The RNN, CRLNet, Informer, and Foundation Model (Single-task) are all trained and evaluated under a full-shot setting, where the target task (time-delay prediction) is directly included during training. These baseline schemes follow the same training configuration as Task II, ensuring a fair and consistent comparison. In contrast, the Foundation Model (1-task), (2-task), and (3-task) variants are trained only on Task I, Task I & Task II, and Task I, Task II & Task III, respectively, without access to the delay prediction task during training. These models are then directly evaluated on the unseen task in a strict zero-shot setting.
@@
As shown in Fig. 8, the proposed foundation model demonstrates strong generalization ability when trained on multiple tasks. Notably, its zero-shot performance on the unseen time-delay prediction task significantly outperforms traditional baselines such as RNN, CRLNet, and Informer. Furthermore, the model trained on all three tasks (i.e., foundation model (3-task)) achieves performance that closely approaches that of the fully supervised foundation model (Single-task), which is directly trained on the target task. In addition, comparing the Foundation Model (1-task), (2-task), and (3-task) variants reveals a clear trend: pretraining on a larger number of diverse tasks consistently improves the model’s zero-shot generalization capability. This confirms the benefit of task diversity in training, and highlights the scalability of the foundation model framework in addressing unseen tasks without requiring retraining or finetuning.

### v2

@@ III-B Model Architecture
Causal Transformer Backbone. We use a causal Transformer to extract features from the input embeddings, consisting of a stack of Transformer layers. Each Transformer layer contains a multi-head causal self-attention mechanism followed by a feed-forward network. For each head $h=\{1,\cdots,H\}$ in the multi-head attention mechanism, we transform $\mathbf{E}^{(m)}$ into a query matrix $\mathbf{Q}^{(m)}_{h}=\mathbf{E}^{(m)}\mathbf{W}^{Q}_{h}$, a key matrix $\mathbf{K}^{(m)}_{h}=\mathbf{E}^{(m)}\mathbf{W}^{K}_{h}$, and a value matrix $\mathbf{V}^{(m)}_{h}=\mathbf{E}^{(m)}\mathbf{W}^{V}_{h}$, where $\mathbf{W}^{Q}_{h},\mathbf{W}^{K}_{h}\in\mathbb{R}^{d_{m}\times d_{k}}$ and $\mathbf{W}^{V}_{h}\in\mathbb{R}^{d_{m}\times d_{v}}$. After that, the $h$-th head after causal attention is given by
@@
where $\mathbf{O}^{(m)}\in\mathbb{R}^{N_{p}\times d_{m}}$ represents the output of multi-head causal attention, and $\mathbf{W}^{O}\in\mathbb{R}^{Hd_{v}\times d_{m}}$ linearly projects these heads to align with the latent dimension of the Transformer. The attention output $\mathbf{O}^{(m)}$ passes through the subsequent feed-forward network and other stacked Transformer layers in the backbone, and is finally converted into a feature matrix $\mathbf{Z}^{(m)}\in\mathbb{R}^{N_{p}\times d_{m}}$.
@@ IV-C1 Baselines and Performance Metric
  • CLRNet: A convolutional long short-term memory (CLSTM) network specifically designed for ISAC scenarios [3]. It combines CNNs for spatial feature extraction with LSTM units to capture temporal dynamics, and is tailored for predictive beamforming via AoA forecasting.
@@ IV-C2 Performance Analysis
We compare the NMSE performance of the foundation models with various baseline schemes across different scenarios, as shown in Fig. 5. As the speeds of the four distinct motion scenarios increase from left to right, a consistent deterioration in performance is observed across all schemes. CLRNet, a model specifically designed for AoA prediction in ISAC, outperforms the naive RNN and Informer networks, both of which were not optimized for ISAC-related tasks. Upon comparing the performance of different schemes, it becomes evident that our proposed foundation models, whether in a single-task or multi-task configuration, significantly outperform the other baseline approaches. Additionally, the single-task model, trained on a specific task, demonstrates a remarkable performance improvement due to its tailored optimization. A more detailed examination of Fig. 5(a) and Fig. 5(b) shows that, in low-speed scenarios (such as Pedestrian and Bicycle), the effect of the prediction interval $\Delta t$ on the performance of all schemes is minimal. In contrast, in high-speed scenarios (such as Vehicle and Rural Vehicle), performance declines as $\Delta t$ increases. This degradation is attributed to the fact that, in high-speed environments, longer prediction intervals result in larger object displacements, which makes it more difficult to predict the AoA accurately. As a result, larger values of $\Delta t$ lead to a more significant reduction in performance.
@@ IV-C3 Downstream Task
Metric RNN CLRNet Informer Foundation Model (Multi-task) Foundation Model (Single-task)
@@ IV-E Zero-Shot and Scaling Analysis
In the previous subsections, we evaluated the zero-shot generalization ability of our foundation model across different configurations within the same task. In this subsection, we take a further step and investigate the model’s zero-shot performance on an unseen task, thereby assessing its capacity for out-of-task generalization. Specifically, we consider a new task: predicting the time-delay in an ISAC system, which corresponds to estimating the physical distance between a user and the BS, i.e. $\nu_{k,t}$ in (8). Structurally, this task resembles previous angle prediction tasks, such as the one described in (II-B). In the following results, we compare the NMSE across time steps for various models to evaluate their performance on the unseen delay prediction task. The RNN, CLRNet, Informer, and Foundation Model (Single-task) are all trained and evaluated under a full-shot setting, where the target task (time-delay prediction) is directly included during training. These baseline schemes follow the same training configuration as Task II, ensuring a fair and consistent comparison. In contrast, the Foundation Model (1-task), (2-task), and (3-task) variants are trained only on Task I, Task I & Task II, and Task I, Task II & Task III, respectively, without access to the delay prediction task during training. These models are then directly evaluated on the unseen task in a strict zero-shot setting.
@@
As shown in Fig. 8, the proposed foundation model demonstrates strong generalization ability when trained on multiple tasks. Notably, its zero-shot performance on the unseen time-delay prediction task significantly outperforms traditional baselines such as RNN, CLRNet, and Informer. Furthermore, the model trained on all three tasks (i.e., foundation model (3-task)) achieves performance that closely approaches that of the fully supervised foundation model (Single-task), which is directly trained on the target task. In addition, comparing the Foundation Model (1-task), (2-task), and (3-task) variants reveals a clear trend: pretraining on a larger number of diverse tasks consistently improves the model’s zero-shot generalization capability. This confirms the benefit of task diversity in training, and highlights the scalability of the foundation model framework in addressing unseen tasks without requiring retraining or finetuning.